# Effortless, Simulation-Efficient Bayesian Inference using Tabular Foundation Models

### v1

@@ Conclusion.
We thank Michael Deistler for suggesting to filter simulations and feedback on the manuscript, and all members of Mackelab for discussions and feedback on the manuscript. This work was funded by the German Research Foundation (DFG) under Germany’s Excellence Strategy – EXC number 2064/1 – 390727645 and SFB 1233 ‘Robust Vision’ (276693517), the German Federal Ministry of Education (Tübingen AI Center) the European Union (ERC, DeepCoMechTome, 101089288), the “Certification and Foundations of Safe Machine Learning Systems in Healthcare” project funded by the Carl Zeiss Foundation and the Boehringer Ingelheim AI & Data Science Fellowship Program. JV and MG are members of the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
Appendix Contents
@@
[…]
Appendix A Software and computational resources
Our implementation of NPE-PF is based on PyTorch [84] and the TabPFN v2 [35] library. We use the SBI library [49] to run the SBI baselines, and hydra [85] for experiment and hyperparameter management. Code to use NPE-PF and reproduce the results is available at https://github.com/mackelab/npe-pf.
We use a mix of Nvidia 2080TI, A100, and H100 GPUs to obtain the results related to NPE-PF. Some results, such as those for the unfiltered NPE-PF variant together with a simulation budget of $10^{5}$ require the use of GPUs with high VRAM due to the large context used. SBI baselines were run on 8 CPU cores, as normalizing flow training and sampling only benefit from GPUs for large sample sizes and dimensionalities.
Appendix B Additional method details
B.1 Filtering to increase effective context size
To allow NPE-PF to make use of more than $10^{4}$ simulations, we filter the simulations based on their relevance to a given observation. In the following, we show that this procedure does not bias the estimation of the posterior and discuss details and possible extensions.
@@ B.1 Filtering to increase effective context size
[…]
@@
[…]
@@
In theory, any function $f$ satisfying these properties can be used to filter for relevant simulations. In this work, we only consider an ABC-like filter, where $f_{{\bm{x}}_{o}}({\bm{x}})=\mathbb{I}(d({\bm{x}},{\bm{x}}_{o})<\epsilon)$ is the indicator function that evaluates to $1$, when some distance $d({\bm{x}},{\bm{x}}_{o})$ between the observation ${\bm{x}}_{o}$ simulation ${\bm{x}}$ is smaller than some threshold $\epsilon>0$, and is 00 otherwise. As described in Sec. 2.3, we use the Euclidean distance in feature-wise standardized observation space and choose $\epsilon$ adaptively to include the $10^{4}$ closest simulations.
The autoregressive manner in which we use TabPFN for density estimation allows for more complex filter designs that consider not only the observation ${\bm{x}}_{o}$, but also already processed parameter dimensions. In particular, it is possible to select parameter-simulation pairs $({\bm{\theta}},{\bm{x}})$ based on their relevance to the parameter dimensions $<$$j$ and observation ${\bm{x}}_{o}$, when using TabPFN to estimate $p(\theta^{j}\mid{\bm{x}}_{o},{\bm{\theta}}^{<j})=q_{\phi}(\theta^{j}\mid{\bm{x}}_{o},{\bm{\theta}}^{<j},{\mathcal{D}}^{<j})$. We do not investigate this approach here and leave the design of more complex (autoregressive) filters to future work.
B.2 Embedding networks for larger feature size
@@
In principle, this end-to-end approach is directly applicable to NPE-PF, but it would require differentiation through the TabPFN model, which can be expensive. We instead aim to obtain lower dimensional embedding, i.e. “summary features” independent from the pre-trained model. This is similar to feature selection schemes usually employed for TabPFN [86, 46].
@@
We refer to this approach as NPE-PF-Infomax and present results on using NPE-PF together with pretrained embedding nets in Sec. D.3..
B.3 Unconditional density estimation
@@
For conditional density estimation (e.g., posterior estimation), filtering provides an effective method to optimize the context given the corresponding condition. However, this strategy does not generalize to unconditional density estimation. One solution is to manually introduce conditional structure. Specifically, we partition the support of the target density into independent clusters and estimate each component separately. Formally, we express the target density as a mixture model
@@
where $p(c_{i})$ is the probability of sampling cluster $c_{i}$, and $p({\bm{x}}\mid c_{i})$ is the cluster-specific density estimated via NPE-PF. This mixture model approach effectively increases the usable context size with the number of clusters. In this work, we apply $k$-means clustering, which has been explored for TabPFN by Xu et al. [75] outside of density estimation. Other methods to introduce conditional structure, such as those proposed by Li et al. [87], represent promising directions for future research. We present results on unconditional density estimation in Sec. D.1.
@@ C.1 SBI Benchmark simulators for amortized and sequential inference
To benchmark NPE-PF and the sequential version TSNPE-PF, we use a set of tasks from the SBI benchmark [24]. This benchmark contains several challenging inference tasks with nonlinear dependencies and multi-modal posteriors. As described in Sec. 3.1, we measure the quality of the inferred posteriors with a classifier-two-sample-test (C2ST). For the classifier, we use a random forest with the default hyperparameters provided by the SBI library [49].
@@
In all experiments, we use the default version of the TabPFN classifier or regressor for (TS)NPE-PF, with no changes to hyperparameters such as the softmax temperature. All runtimes for NPE-PF (Fig. 2b) were obtained using an Nvidia A100 GPU, where possible. For the unfiltered variant of NPE-PF, an H100 GPU was used for the large context containing $10^{5}$ simulations.
@@ C.2 Misspecification simulator
For the experiments, we ran 100 simulations for three seeds over all combinations of $(\mu_{m},\tau_{m},\lambda)$. NPE and NPE-PF are trained to predict posterior samples. These samples are compared to the reference posterior distribution using C2ST. This reference posterior distribution is defined with respect to the well-specified distribution. Thus, this task is designed to evaluate the robustness of the model under prior or likelihood misspecification.
@@ C.3 Single-compartment Hodgkin-Huxley simulator
To perform inference on the observations from the Allen cell types database [56], we use a single-compartment Hodkin-Huxley model following the one proposed in Pospischil et al. [54]. This model has four types of conductances (sodium, delayed-rectifier potassium, slow voltage-dependent potassium, and leak) and a total of eight parameters. Following Gonçalves et al. [55], who previously used this model for simulation-based inference, we compute seven summary statistics on the simulated and observed action potentials. These are spike count, mean, and standard deviation of the resting potential, and the first four voltage moments (mean, standard deviation, skew, and kurtosis). We also use the same uniform prior as in Gonçalves et al. [55]. For sequential inference with TSNPE and TSNPE-PF, we use a total of $10^{3}$ or $10^{4}$ simulations, equally divided over five rounds. For both methods, we sample from the truncated proposal using rejection sampling taking into account the $1-\varepsilon$ highest density region with $\varepsilon=10^{-3}$.
Here, we provide posterior predictive samples for all 10 observations from the Allen cell types database using TSNPE-PF (Fig. D-8) and TSNPE (Fig. D-9), complementing the results presented in the main text (Fig. 4). Importantly both inference and the computation of the average predictive distance are performed on the summary statistics and not directly on the simulations or observations. Thus, features beyond those encoded by the summary statistics are not captured. For TSNPE-PF, the posterior predictives closely match with the observations in most cases, with a few exceptions such as obs. 3 and obs. 6. In contrast, flow-based TSNPE yields posterior predictives that are less similar to the observations, with a few exceptions such as obs. 1 and obs. 8. These qualitative results further support the quantitative results reported in the main text, demonstrating that TNSPE-PF performs more reliable inference with fewer simulations.
@@ C.4 Pyloric Network Model
A key challenge in this setting is the extreme sparsity of valid simulations. In particular, 99% of parameter samples from the prior yield implausible voltage traces, preventing the computation of summary statistics. For this reason, previous work has relied on millions of simulations for an amortized NPE-based posterior approximation (18 million in Gonçalves et al. [55]). This could later be reduced to 9 million by restricting the prior to valid simulations with an additional classifier [62]. Moreover, several sequential algorithms have been developed to tackle this problem. First, SNVI was proposed, which learns both a likelihood and a variational approximation of the posterior. Notably, the likelihood model was only based on “valid” simulations and corrected by a classifier that predicts whether parameters are valid. This approach made it possible to produce good posterior samples with $3.5\cdot 10^{5}$ simulations [18]. This number was reduced even further to only $1.5\cdot 10^{5}$ simulations by S-UNLE, which made use of energy-based models and Langevin dynamic MCMC [63]. Later, this approach was extended to SNPE approximations, specifically using TSNPE [34].
@@
Similarly, we want to use only valid simulations in the context of NPE-PF and therefore adaptively truncate the prior using a TabPFN classifier. This approach is based on the intuition that the posterior should place little or no probability mass on regions in parameter space that are likely to lead to invalid simulations [62, 34, 18]. Specifically, as simulations accumulate over rounds, we randomly fill the context of the TabPFN classifier with up to $5000$ valid and invalid simulations. The prior is then truncated to $\tilde{p}_{\text{val}}({\bm{\theta}})\propto\mathbb{I}(P(\text{valid}\mid{\bm{\theta}})>c)p({\bm{\theta}})$, where we choose $c=0.3$. We then use TSNPE-PF together with an approximate importance resampling scheme instead of rejection sampling for efficiency. Specifically, we oversample by a factor of $K=10$ and obtain $10^{4}$ samples, which are then resampled to $10^{3}$ samples based on their importance weight $w({\bm{\theta}})=\tilde{p}_{\text{val}}({\bm{\theta}})/q_{\phi}({\bm{\theta}}\mid{\bm{x}}_{o})$. For evaluation, we exactly follow Glaser et al. [63] by computing the percentage of valid simulations as well as the energy score in each round. Note, however, that the overall experimental setting is not perfectly identical as we use a faster implementation of the pyloric network simulator based on Jaxley [89].
@@
Previous sequential methods started with an initial simulation budget of $5\cdot 10^{4}$ prior simulations to acquire a set of $\sim 500$ valid simulations. Due to the effectiveness of NPE-PF on small simulation budgets, we start with only $5000$ simulations, thus acquiring only $\sim 50$ valid simulations in the first round. From there, we acquire $1000$ new simulations in each round, updating both NPE-PF and the restricted classifier with the additional data. We stop after 45 rounds for a total of $5\cdot 10^{4}$ simulations. Despite this limited budget, TSNPE-PF quickly converges to a posterior that produces a high number of valid simulations that closely match the observed measurement. After the final round, we obtain a valid percentage of $96.83\%$ and an energy score of $0.0899$ using only $5\cdot 10^{4}$ simulations—the starting point of the other methods—and still surpass their final performance (Fig. 5). The evaluation presented in the main text only examines the fidelity of the posterior predictives. Therefore, we include a visualization of the full posterior distribution after the final round (Fig. D-10), which shows broad marginal distributions and similar features as the posteriors reported in previous work [55, 18, 63, 34].
@@
In terms of runtime, most of the computation time is spent on sampling from the restricted prior and running simulations. During the first few rounds, the process is faster (about 2 minutes) due to the smaller context size in TSNPE-PF. Once the context limit is reached, the time to propose the next 1000 parameters stabilizes at about 7 minutes, using an NVIDIA H100 GPU. Notably, the performance of previous state-of-the-art methods on this task is surpassed after only 10 rounds—an hour of runtime (including simulation time)—which is substantially faster than any previously proposed methods [63, 18].

### v2

@@ 4 Discussion
We thank Michael Deistler and Katharina Eggensperger for feedback on the manuscript, specifically Michael Deistler for suggesting to filter simulations and Jonas Beck for the implementation of the pyloric simulator. We thank all members of the Mackelab for discussions and feedback on the manuscript. This work was funded by the German Research Foundation (DFG) under Germany’s Excellence Strategy – EXC number 2064/1 – 390727645 and SFB 1233 ‘Robust Vision’ (276693517), the European Union (ERC, DeepCoMechTome, 101089288), the “Certification and Foundations of Safe Machine Learning Systems in Healthcare” project funded by the Carl Zeiss Foundation and the Boehringer Ingelheim AI & Data Science Fellowship Program. JV and MG are members of the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
Appendix Table of Contents
Appendix A Software and computational resources
Our implementation of NPE-PFN is based on PyTorch [86] and the TabPFN [38] library. We use the SBI library [54] to run the SBI baselines, and hydra [87] for experiment and hyperparameter management. Code to use NPE-PFN and reproduce the results is available at https://github.com/mackelab/npe-pfn.
We use a mix of Nvidia 2080TI, A100, and H100 GPUs to obtain the results related to NPE-PFN. Some results, such as those for the unfiltered NPE-PFN variant together with a simulation budget of $10^{5}$ require the use of GPUs with high VRAM due to the large context used. SBI baselines were run on 8 CPU cores, as normalizing flow training and sampling only benefit from GPUs for large sample sizes and dimensionalities.
Appendix B Additional method details
B.1 Pseudocode: NPE-PFN and TSNPE-PFN
In Alg. 1 and Alg. 2, we provide pseudocode for NPE-PFN and TSNPE-PFN, respectively. Note that NPE-PFN is training-free. In practice, this means that training consists of storing the training data ${\mathcal{D}}=\{({\bm{\theta}}_{i},{\bm{x}}_{i})\}_{i=1}^{N}$ for in-context processing during inference.
Algorithm 1 NPE-PFN
1:Observation ${\bm{x}}_{o}$, simulations ${\mathcal{D}}_{\mathrm{full}}=\{({\bm{\theta}}_{i},{\bm{x}}_{i})\}_{i=1}^{N}$, TabPFN regressor $q_{\psi}^{\text{reg}}(\cdot\mid\cdot)$, filter size $N_{\mathrm{filter}}$
2:if $N_{\mathrm{filter}}<N$ then $\triangleright$ if filtering is active
3:  $d_{i}=d({\bm{x}}_{o},{\bm{x}}_{i})$ $\triangleright$ compute filter distances
4:  ${\mathcal{D}}\leftarrow\{({\bm{\theta}}_{i},{\bm{x}}_{i}):i\in\mathrm{argtop}_{N_{\mathrm{filter}}}(d_{i})\}$ $\triangleright$ select $N_{\mathrm{filter}}$ closest simulations
5:else
6:  ${\mathcal{D}}\leftarrow{\mathcal{D}}_{\mathrm{full}}$
7:end if
8:for $j=1,\dots,d_{\bm{\theta}}$ do $\triangleright$ for all parameter dimension $d_{\bm{\theta}}$
9:  ${\mathcal{D}}^{<j}=\{\theta_{i}^{j},[{\bm{\theta}}_{i}^{<j},{\bm{x}}_{i}]\}$ $\triangleright$ build context data set
10:  $\theta^{j}\sim q_{\psi}^{\text{reg}}(\theta^{j}\mid{\bm{\theta}}^{<j},{\bm{x}}_{o},{\mathcal{D}}^{<j})$ $\triangleright$ sample one parameter dimension from TabPFN
11:end for
12:return ${\bm{\theta}}=[\theta_{1},\theta_{2},\ldots,\theta_{d_{\theta}}]^{\top}$ $\triangleright$ posterior sample
Algorithm 2 TSNPE-PFN with rejection sampling
1:Observation ${\bm{x}}_{o}$, simulator $p({\bm{x}}\mid{\bm{\theta}})$, prior $p({\bm{\theta}})$, TabPFN classifier $q_{\psi}^{\text{cls}}(\cdot\mid\cdot)$, TabPFN regressor $q_{\psi}^{\text{reg}}(\cdot\mid\cdot)$, number of rounds $R$, number of posterior samples for ratio $M$, simulations per round $N_{r}$, truncation fraction $\alpha$
2:— Initialize dataset —
3:Draw $N_{r}$ prior samples ${\bm{\theta}}_{i}\sim p({\bm{\theta}})$
4:Simulate samples ${\bm{x}}_{i}\sim p({\bm{x}}\mid{\bm{\theta}}_{i})$
5:Initialize ${\mathcal{D}}=\{({\bm{\theta}}_{i},{\bm{x}}_{i})\}_{i=1}^{N_{r}}$
6:
7:for $r=1,\dots,R$ do
8:  — Construct fast ratio-based density estimator —
9:  Draw $M$ posterior samples ${\bm{\theta}}_{j}^{\text{post}}\sim q({\bm{\theta}}\mid{\bm{x}}_{o})$ using Alg. 1 with $\mathcal{D}$ as context
10:  Assign class $y=1$ to ${\bm{\theta}}_{j}^{\text{post}}$ and $y=0$ to ${\bm{\theta}}_{j}^{\text{uni}}\sim\mathcal{U}({\bm{\theta}}_{\min},{\bm{\theta}}_{\max})$
11:  Define classifier dataset ${\mathcal{D}}_{\text{cls}}=\{({\bm{\theta}}_{j},y_{j})\}_{j=1}^{2M}$ combining ${\bm{\theta}}_{j}^{\text{post}}$ and ${\bm{\theta}}_{j}^{\text{uni}}$
12:  With ${\mathcal{D}}_{\text{cls}}$ as context, $q_{\psi}^{\text{cls}}$ can be used to estimate posterior density:
$q_{\text{cls}}({\bm{\theta}}\mid{\bm{x}}_{o})=\frac{q_{\psi}^{\text{cls}}(y=1\mid{\bm{\theta}},{\mathcal{D}}_{\text{cls}})}{1-q_{\psi}^{\text{cls}}(y=1\mid{\bm{\theta}},{\mathcal{D}}_{\text{cls}})}.$
13:
14:  — Estimate high-density region (HDR) —
15:  $\text{qs}=\{q_{\text{cls}}({\bm{\theta}}_{j}^{\text{post}}|\mathbf{x}_{o})\ \ \text{for}\ j=1,\dots,M\}\triangleright$ evaluate posterior densities
@@ B.1 Pseudocode: NPE-PFN and TSNPE-PFN
17:
18:  — Rejection sample from truncated prior —
19:  $D_{r}=\emptyset$
20:  while $|D_{r}|<N_{r}$ do
21:   ${\bm{\theta}}\sim p({\bm{\theta}})$ $\triangleright$ sample from prior
22:   if $q_{\text{cls}}({\bm{\theta}}\mid\mathbf{x}_{o})\geq k_{\alpha}$ then $\triangleright$ accept if in HDR
23:     $\mathbf{x}\sim p(\mathbf{x}\mid{\bm{\theta}})$
24:     $D_{r}\leftarrow D_{r}\cup\{({\bm{\theta}},\mathbf{x})\}$
25:   end if
26:  end while
27:  $\mathcal{D}=\mathcal{D}\cup\mathcal{D}_{r}$
28:end for
29:
30:— Sample final samples —
31:Draw posterior samples ${\bm{\theta}}_{i}\sim q({\bm{\theta}}\mid{\bm{x}}_{o})$ using Alg. 1 with $\mathcal{D}$} as context
32:return $\{{\bm{\theta}}_{i}\}_{i}$
B.2 Filtering to increase effective context size
To allow NPE-PFN to make use of more than $10^{4}$ simulations, we filter the simulations based on their relevance to a given observation. In the following, we show that this procedure does not bias the estimation of the posterior and discuss details and possible extensions.
@@ B.2 Filtering to increase effective context size
$p({\bm{\theta}}\mid{\bm{x}})=\frac{p({\bm{x}}\mid{\bm{\theta}})p({\bm{\theta}})}{p({\bm{x}})}=\frac{p({\bm{x}},{\bm{\theta}})}{\int p({\bm{x}},{\bm{\theta}}^{\prime})\,d{\bm{\theta}}^{\prime}}.$ (B-1)
@@
$\tilde{p}({\bm{\theta}}\mid{\bm{x}})=\frac{\tilde{p}({\bm{x}},{\bm{\theta}})}{\int\tilde{p}({\bm{x}},{\bm{\theta}}^{\prime})\,d{\bm{\theta}}^{\prime}}=\frac{f_{{\bm{x}}_{o}}({\bm{x}})p({\bm{x}},{\bm{\theta}})}{\int f_{{\bm{x}}_{o}}({\bm{x}})p({\bm{x}},{\bm{\theta}}^{\prime})\,d{\bm{\theta}}^{\prime}}=p({\bm{\theta}}\mid{\bm{x}}),$ (B-2)
@@
In theory, any function $f$ satisfying these properties can be used to filter for relevant simulations. In this work, we only consider an ABC-like filter, where $f_{{\bm{x}}_{o}}({\bm{x}})=\mathbb{I}(d({\bm{x}},{\bm{x}}_{o})<\epsilon)$ is the indicator function that evaluates to $1$, when some distance $d({\bm{x}},{\bm{x}}_{o})$ between the observation ${\bm{x}}_{o}$ simulation $\bm{x}$} is smaller than some threshold $\epsilon>0$, and is 0 otherwise. As described in Sec. 2.3, we use the Euclidean distance in feature-wise standardized observation space and choose $\epsilon$ adaptively to include the $10^{4}$ closest simulations.
The autoregressive manner in which we use TabPFN for density estimation allows for more complex filter designs that consider not only the observation ${\bm{x}}_{o}$, but also already processed parameter dimensions. In particular, it is possible to select parameter-simulation pairs $({\bm{\theta}},{\bm{x}})$ based on their relevance to the parameter dimensions <<$j$ and observation ${\bm{x}}_{o}$, when using TabPFN to estimate $p(\theta^{j}\mid{\bm{x}}_{o},{\bm{\theta}}^{<j})=q_{\psi}(\theta^{j}\mid{\bm{x}}_{o},{\bm{\theta}}^{<j},{\mathcal{D}}^{<j})$. We do not investigate this approach here and leave the design of more complex (autoregressive) filters to future work.
B.3 Embedding networks for high-dimensional data
@@
In principle, this end-to-end approach is directly applicable to NPE-PFN, but it would require differentiation through the TabPFN model, which can be expensive. We instead aim to obtain lower dimensional embedding, i.e., “summary features” independent from the pre-trained model. This idea is similar to feature selection schemes usually employed for TabPFN [88, 51].
@@
We refer to this approach as NPE-PFN-Infomax and present results on using NPE-PFN together with pre-trained embedding nets in Sec. D.4.
B.4 Unconditional density estimation
@@
For conditional density estimation (e.g., posterior estimation), filtering provides an effective method to optimize the context given the corresponding condition. However, this strategy does not generalize to unconditional density estimation. One solution is to manually introduce a conditional structure. Specifically, we partition the support of the target density into independent clusters and estimate each component separately. Formally, we express the target density as a mixture model
@@
where $p(c_{i})$ is the probability of sampling cluster $c_{i}$, and $p({\bm{x}}\mid c_{i})$ is the cluster-specific density estimated via NPE-PFN. This mixture model approach effectively increases the usable context size with the number of clusters. In this work, we apply $k$-means clustering, which has been explored for TabPFN by Xu et al. [78] outside of density estimation. Other methods to introduce conditional structure, such as those proposed by Li et al. [89], represent promising directions for future research. We present results on unconditional density estimation in Sec. D.9.
@@ C.1 SBI benchmark simulators for amortized and sequential inference
Table C-1: Overview on the SBI benchmark tasks. For details on the specification of the prior and the exact simulator equations, we refer to Lueckmann et al. [27, Appendix T].
Task Name $\dim{\bm{\theta}}$ $\dim{\bm{x}}$ Notes
Gaussian Linear 10 10 Mean inference in 10-D Gaussian with fixed covariance
Gaussian Mixture 2 2 Common mean of two 2-D Gaussians with varying covariance
Two Moons 2 2 Bimodal, two moon-shaped posterior structure
SLCP 5 8 Simple Gaussian likelihood, complex posterior
Bernoulli GLM 10 10 GLM with Bernoulli observations, smoothness prior
SIR 2 10 Epidemic SIR model, inference of contact and recovery rates
Lotka-Volterra 4 20 Predator-prey dynamics, species interaction parameters
To benchmark NPE-PFN and the sequential version TSNPE-PFN, we use a set of tasks from the SBI benchmark [27]. This benchmark contains several challenging inference tasks with nonlinear dependencies and multi-modal posteriors. We provide an overview of the dimensionality of parameters and observations for all benchmark tasks (Tab. C-1). As described in Sec. 3.1, we measure the quality of the inferred posteriors with a classifier-two-sample-test (C2ST). For the classifier, we use a random forest with the default hyperparameters provided by the SBI library [54].
@@
In all experiments, we use the default version of the TabPFN classifier or regressor for (TS)NPE-PFN, with no changes to hyperparameters such as the softmax temperature. All runtimes for NPE-PFN (Fig. 2b) were obtained using an Nvidia A100 GPU, where possible. For the unfiltered variant of NPE-PFN, an H100 GPU was used for the large context containing $10^{5}$ simulations.
@@ C.2 Misspecification simulator
For the experiments, we ran 100 simulations for three seeds over all combinations of $(\mu_{m},\tau_{m},\lambda)$. NPE and NPE-PFN are trained to predict posterior samples. These samples are compared to the reference posterior distribution using C2ST. This reference posterior distribution is defined with respect to the well-specified distribution. Thus, this task is designed to evaluate the robustness of the model under prior or likelihood misspecification.
@@ C.3 Single-compartment Hodgkin-Huxley simulator
To perform inference on the observations from the Allen cell types database [61], we use a single-compartment Hodkin-Huxley model following the one proposed in Pospischil et al. [59]. This model has four types of conductances (sodium, delayed-rectifier potassium, slow voltage-dependent potassium, and leak) and a total of eight parameters. Following Gonçalves et al. [60], who previously used this model for simulation-based inference, we compute seven summary statistics on the simulated and observed action potentials. These are spike count, mean, and standard deviation of the resting potential, and the first four voltage moments (mean, standard deviation, skew, and kurtosis). We also use the same uniform prior as in Gonçalves et al. [60]. For sequential inference with TSNPE and TSNPE-PFN, we use a total of $10^{3}$ or $10^{4}$ simulations, equally divided over five rounds. For both methods, we sample from the truncated proposal using rejection sampling, taking into account the $1-\varepsilon$ highest density region with $\varepsilon=10^{-3}$.
Here, we provide posterior predictive samples for all 10 observations from the Allen cell types database using TSNPE-PFN (Fig. D-11) and TSNPE (Fig. D-12), complementing the results presented in the main text (Fig. 4). Importantly, both inference and the computation of the average predictive distance are performed on the summary statistics and not directly on the simulations or observations. Thus, features beyond those encoded by the summary statistics are not captured. For TSNPE-PFN, the posterior predictives closely match with the observations in most cases, with a few exceptions such as obs. 3 and obs. 6. In contrast, flow-based TSNPE yields posterior predictives that are less similar to the observations, with a few exceptions such as obs. 1 and obs. 8. These qualitative results further support the quantitative results reported in the main text, demonstrating that TNSPE-PF performs more reliable inference with fewer simulations.
@@ C.4 Pyloric Network Model
A key challenge in this setting is the extreme sparsity of valid simulations. In particular, 99% of parameter samples from the prior yield implausible voltage traces, preventing the computation of summary statistics. For this reason, previous work has relied on millions of simulations for an amortized NPE-based posterior approximation (18 million in Gonçalves et al. [60]). The simulation count could later be reduced to 9 million by restricting the prior to valid simulations with an additional classifier [67]. Moreover, several sequential algorithms have been developed to tackle this problem. First, SNVI was proposed, which learns both a likelihood and a variational approximation of the posterior. Notably, the likelihood model was only based on “valid” simulations and corrected by a classifier that predicts whether parameters are valid. This approach made it possible to produce good posterior samples with $3.5\cdot 10^{5}$ simulations [18]. This number was reduced even further to only $1.5\cdot 10^{5}$ simulations by S-UNLE, which made use of energy-based models and Langevin dynamic MCMC [68]. Later, this approach was extended to SNPE approximations, specifically using TSNPE [37].
@@
Similarly, we want to use only valid simulations in the context of NPE-PFN and therefore adaptively truncate the prior using a TabPFN classifier. This approach is based on the intuition that the posterior should place little or no probability mass on regions in parameter space that are likely to lead to invalid simulations [67, 37, 18]. Specifically, as simulations accumulate over rounds, we randomly fill the context of the TabPFN classifier with up to $5\cdot 10^{3}$ valid and invalid simulations. The prior is then truncated to $\tilde{p}_{\text{val}}({\bm{\theta}})\propto\mathbb{I}(P(\text{valid}\mid{\bm{\theta}})>c)p({\bm{\theta}})$, where we choose $c=0.3$. We then use TSNPE-PFN together with an approximate importance resampling scheme instead of rejection sampling for efficiency. Specifically, we oversample by a factor of $K=10$ and obtain $10^{4}$ samples, which are then resampled to $10^{3}$ samples based on their importance weight $w({\bm{\theta}})=\tilde{p}_{\text{val}}({\bm{\theta}})/q_{\phi}({\bm{\theta}}\mid{\bm{x}}_{o})$. For evaluation, we exactly follow Glaser et al. [68] by computing the percentage of valid simulations as well as the energy score in each round. Note, however, that the overall experimental setting is not perfectly identical as we use a faster implementation of the pyloric network simulator based on Jaxley [91].
@@
Previous sequential methods started with an initial simulation budget of $5\cdot 10^{4}$ prior simulations to acquire a set of $\sim 500$ valid simulations. Due to the effectiveness of NPE-PFN on small simulation budgets, we start with only $5\cdot 10^{3}$ simulations, thus acquiring only $\sim 50$ valid simulations in the first round. From there, we acquire $10^{3}$ new simulations in each round, updating both NPE-PFN and the restricted classifier with the additional data. We stop after 45 rounds for a total of $5\cdot 10^{4}$ simulations. Despite this limited budget, TSNPE-PFN quickly converges to a posterior that produces a high number of valid simulations that closely match the observed measurement. After the final round, we obtain a valid percentage of $96.83\%$ and an energy score of $0.0899$ using only $5\cdot 10^{4}$ simulations—the starting point of the other methods—and still surpass their final performance (Fig. 5). The evaluation presented in the main text only examines the fidelity of the posterior predictives. Therefore, we include a visualization of the full posterior distribution after the final round (Fig. D-13), which shows broad marginal distributions and similar features as the posteriors reported in previous work [60, 18, 68, 37].
@@
In terms of runtime, most of the computation time is spent on sampling from the restricted prior and running simulations. During the first few rounds, the process is faster (about 2 minutes) due to the smaller context size in TSNPE-PFN. Once the context limit is reached, the time to propose the next $10^{3}$ parameters stabilizes at about 7 minutes, using an Nvidia H100 GPU. Notably, the performance of previous state-of-the-art methods on this task is surpassed after only 10 rounds—an hour of runtime (including simulation time)—which is substantially faster than any previously proposed methods [68, 18].
We choose the default order as the autoregressive order for sampling. The total number of possible permutations in this task is greater than $10^{33}$, thus finding the optimal permutation in this space is highly challenging. To gain some insight, we perform an ablation study on 50 distinct random permutations using the context dataset identified by the final‑round posterior. We compute the presented metrics with $10^{3}$ simulations. We found that the energy score varies between 0.081 - 0.094 and the valid rates between $96\%$ - $98\%$ ($5\%$ and $95\%$ quantile each). These ranges match closely with the results reported in Fig. 5 and demonstrate that there is no significant effect for the autoregressive order in this experiment. Appendix Sec. D.8 reports the same conclusion across all benchmark tasks.